A company has just paid a $2 dividend per share. The dividend is expected to grow at a constant rate of 9% per year for 3 years. After that, the dividend is expected to keep growing at a constant rate of 3% per year. The required return is 9%. What will the stock price be in 2 years?
Answer #1

Last Dividend, D0 = $2.00

Growth rate for next 3 years is 9% and a constant growth rate (g) of 3% thereafter

D1 = $2.0000 * 1.09 = $2.1800
D2 = $2.1800 * 1.09 = $2.3762
D3 = $2.3762 * 1.09 = $2.5901
D4 = $2.5901 * 1.03 = $2.6678

Required Return, rs = 9%

P3 = D4 / (rs - g)
P3 = $2.6678 / (0.09 - 0.03)
P3 = $2.6678 / 0.06
P3 = $44.4633

P2 = D3 / (1 + rs) + P3 / (1 + rs)
P2 = $2.5901 / 1.09 + $44.4633 / 1.09
P2 = $43.17

Therefore, stock price in 2 years will be $43.17